Show how you solve the following two equations:
6x = 33
2/3x=8/15

Answer by Alan3354(69443)

Show how you solve the following two equations:
6x = 33
Divide by 6
--------
2/3x=8/15
If it's (2/3)x = 8/15
divide by 2/3
============
If it's 2/(3x) = 8/15
Cross multiply, then divide by 3